OpenStudy (superhelp101):

Which system of equations matches the graph below? look for the answer choices below

OpenStudy (anonymous):

graph?

OpenStudy (superhelp101):

@mathmate or aum :)

OpenStudy (aum):

The graph shows a straight line with a negative slope when x <= 1. And a parabola when x > 1. Therefore, ....

OpenStudy (aum):

\( \le \) is read as "less than or equal to" \( \ge \) is read as "greater than or equal to"

OpenStudy (aum):

You may want to change the wordings in your answer choices a, b and d.

OpenStudy (superhelp101):

why?

OpenStudy (aum):

What does "x less than or greater to 1" mean?

OpenStudy (superhelp101):

i'm not sure let me check

OpenStudy (superhelp101):

I will re edit the question :)

OpenStudy (superhelp101):

y = x2 if x < 1 y = -x if x less than or equal to 1 y = -x if x less than or equal to 1 y = x2 if x > 1 y = x2 if x > 1 y = x if x greater than or equal to -1 y = x if x less than or equal to 1 y = -x2 if x > 1

OpenStudy (superhelp101):

each two lines are one answer choice

OpenStudy (aum):

You have two parts to the graph. A straight line on the left and part of a parabola on the right. The domain of the straight line is x≤1. Notice how it is ≤ and not \(<\). That is because at x = 1, the straight line has a closed circle indicating x = 1 is part of the straight line. But in the parabola, at x = 1 you see an open circle indicating that x = 1 is not part of the parabola. Now the straight line has a negative slope and it is a 45 degree straight line, passing through the origin. Therefore, the equation of the straight line is: y = -1 for x≤1.

OpenStudy (aum):

I meant y = -x for x≤1.

OpenStudy (aum):

The parabola is \(y = x^2\) for \(x \gt 1\)

OpenStudy (aum):

Yes, B.

OpenStudy (superhelp101):

oh ok thx! :)

OpenStudy (aum):

yw.